### 09 Feb 2024 v3

# Cyber-Twin: Digital Twin-boosted Autonomous Attack Detection for Vehicular Ad-Hoc Networks

Abstract
The rapid evolution of Vehicular Ad-hoc NETworks (VANETs) has ushered in a transformative era for intelligent transportation systems (ITS), significantly enhancing road safety and vehicular communication. However, the intricate and dynamic nature of VANETs presents formidable challenges, particularly in vehicle-to-infrastructure (V2I) communications. Roadside Units (RSUs), integral components of VANETs, are increasingly susceptible to cyberattacks, such as jamming and distributed denial-of-service (DDoS) attacks. These vulnerabilities pose grave risks to road safety, potentially leading to traffic congestion and vehicle malfunctions. Current approaches often struggle to effectively merge digital twin technology with Artificial Intelligence (AI) models to boost security and sustainability. In response to this gap, our study introduces an innovative cyber-twin framework tailored to enhance the security of RSUs in VANETs. This framework uniquely combines digital twin technology with cutting-edge AI to offer a real-time, dynamic representation of RSUs. This allows for detailed monitoring and efficient detection of threats, significantly strengthening RSU security in VANETs. Moreover, our framework makes a notable contribution to eco-friendly communication by improving the computational efficiency of RSUs, leading to increased energy efficiency and extended hardware durability. Our results show a considerable enhancement in resource management and attack detection, surpassing the performance of existing solutions. In particular, the cyber-twin framework showed a substantial reduction in RSU load and an optimal balance between resource consumption and high attack detection efficiency, with a defined twinning rate range of seventy-six to ninety per cent. These advancements underscore our commitment to developing sustainable, secure, and resilient vehicular communication systems for the future of smart cities.
Index Terms:
VANETs, ITS, Digital Twin, Cybersecurity in Transportation, Green Communication.
I Introduction
In recent years, the advent of Vehicular Ad-hoc NETworks (VANETs) has marked a significant milestone in the evolution of intelligent transportation systems (ITS). These networks, which seamlessly integrate mobile ad hoc networks (MANET), the Internet of Things (IoT), and intelligent transport systems, are at the forefront of revolutionizing road safety and vehicular communication. The critical role of VANETs in reducing traffic accidents and enhancing road safety cannot be overstated, given the alarming statistics from the World Health Organization highlighting the high incidence of road traffic deaths globally [1].
However, the dynamic and complex nature of VANETs introduces significant challenges, particularly in vehicle-to-infrastructure communications [2]. Road Side Units (RSUs), as critical components of VANETs, are often the targets of cyberattacks, including but not limited to jamming and distributed denial-of-service (DDoS) attacks, which can severely disrupt vehicular communication and pose significant risks to road safety, leading to potential traffic congestion, vehicle malfunctions, and even accidents [3]. This vulnerability underscores the urgency for innovative solutions to safeguard these systems against such threats. The urgent need to address these vulnerabilities is underscored by the rapid evolution of vehicular technology and the increasing prevalence of connected and autonomous vehicles in modern urban environments.
Figure 1: The cyber-twin architecture.
In addressing these challenges, our work focuses primarily on vehicle-to-infrastructure communication, explicitly targeting the security of RSUs. We introduce a novel cyber-twin framework that employs digital twin technology and an advanced Artificial Intelligence (AI) model. The digital twin concept, a digital replica of a physical system, is central to our approach [4]. By creating a digital twin of RSUs, we achieve a real-time, dynamic representation of the physical units, which allows for the meticulous monitoring and analysis of network interactions. This capability is crucial in detecting and mitigating the aforementioned cyber threats, ensuring the integrity and reliability of vehicle-to-infrastructure communications [5] Furthermore, the significance of our solution lies not only in its security aspects but also in its contribution to green communication. By optimizing the efficiency of RSUs and reducing their computational demands, our framework aligns with the growing emphasis on sustainable and eco-friendly technologies in smart cities. This dual focus on security and sustainability positions our work at the forefront of advancing intelligent transportation systems. Moreover, this reduction in energy consumption and prolongation of hardware lifespans underscores our commitment to developing environmentally conscious technologies for smart cities. The key contributions of this paper include:
  • We propose a cyber-twin framework tailored explicitly for securing vehicle-to-infrastructure communication in VANETs, utilizing an AI model for efficient threat detection.
  • By integrating digital twin technology, our approach provides a novel and effective method for real-time monitoring, management of vehicular networks, and analysis of RSU performance and security.
  • The cyber-twin framework significantly reduces the computational load on RSUs, contributing to the longevity of hardware and overall energy efficiency.
  • Our solution has significant advancements in green communication practices, contributing to the sustainability of intelligent transportation systems.
This paper presents a pioneering approach to enhancing the security and efficiency of vehicle-to-infrastructure communication in VANETs while simultaneously addressing the critical aspects of road safety and environmental sustainability. We aim to set a new standard in developing intelligent and secure transportation systems through the novel application of digital twin technology and AI models. In essence, our work addresses a critical gap in vehicular network security. It paves the way for more sustainable and intelligent transportation systems, marking a significant stride in the journey towards safer and greener smart cities. This work proceeds with an overview of relevant literature in Section II, then delves into the proposed methodology in Section III, and evaluates its performance in Section IV. The paper is drawn to a close with concluding remarks in Section V.
II Related Work
Recent research in VANETs and their integration into ITS has significantly focused on improving road safety and efficiency. This section reflects on key advancements in this domain, setting the stage for our novel cyber-twin framework.
Alhaidari et al. introduced a simulation method to create the VANET DDoS dataset, encompassing critical VANET features like traffic density and node mobility. However, it remains inaccessible to the public [6]. Bangui et al. tackled VANETs’ vulnerability to malicious jamming, proposing a data prioritization model enhancing Big Data Analytics’ efficiency in jamming detection, a leap forward in real-time anti-jamming applications [7]. Mokhtar et al. surveyed VANETs’ security, highlighting the challenges due to their mobile, infrastructure-less nature [2]. They emphasized the network’s inherent security vulnerabilities, paving the way for innovative solutions. Another notable study introduced a machine learning-based system for real-time identification of malicious nodes in VANETs, achieving high accuracy with Random Forest (RF) and Gradient-boosted Trees (GBT) models [8].
Figure 2: The state diagram
Haydari et al. developed advanced machine learning methods for intrusion detection in VANETs, targeting sophisticated attacks like DDoS [9]. Their focus on RSUs brought forward a non-parametric IDS, though its effectiveness against sporadic attacks remains a potential area for improvement. Zhou et al. concentrated on enhancing VANET security within ITS for 6G systems, introducing a framework combining identity-based encryption and deep learning, primarily addressing DDoS attack scenarios [10]. Alrehan et al. reviewed VANET architectures, advocating for security solutions uniquely tailored to VANET characteristics, diverging from conventional methods [11]. Anyanwu et al. focused on software-defined network (SDN) integrated VANETs’ vulnerability to DDoS attacks led to an Intrusion Detection Model utilizing a support vector machine classifier, a significant step in vehicle security [12]. A novel self-powered, fog-based VANET infrastructure was also proposed, focusing on sustainable and secure communication, including innovative power management and fog clustering strategies for enhanced network resilience [13].
Within VANET research, the application of digital twin technology remains a relatively underexplored area. Among the few, Arya et al. utilized it for identifying malicious nodes, employing machine learning for practical traffic analysis [14]. Ak et al. developed T6CONF, a new framework that enhances IoT communication in smart cities, thereby improving waste management predictions and achieving net-zero waste objectives [15]. Additionally, Hu et al. explored the Internet of Vehicles (IoV), forming a digital twin connection between physical and virtual vehicles for real-time traffic prediction, highlighting the potential and challenges in handling large, sparse data sets [16].
These studies underscore the untapped potential of VANETs for road safety and network security. However, there is a noticeable gap in effectively combining digital twin technology with AI models for enhanced security and environmental sustainability in VANETs. Our proposed cyber-twin framework aims to bridge this gap, presenting a novel approach to secure vehicle-to-infrastructure communication and advancing green communication in smart city infrastructures.
III Proposed Cyber-twin System Model for Vehicle-to-Infrastructure Security
Our proposed system is designed to foster a resilient and robust communication framework for vehicles in 6G smart cities, leveraging digital twin technology to optimize the computational efficiency of roadside units (RSUs). This enhances the reliability and stability of the RSUs and significantly contributes to green communication initiatives. Some potential benefits of our system can be:
  • Enhanced energy efficiency in RSUs, achieved through reduced computational demands.
  • Extension of hardware lifespan, resulting in economic benefits due to decreased replacement and maintenance needs.
  • Greater adaptability of RSUs to emerging technologies, eliminating the need for comprehensive hardware overhauls.
Our system adopts a holistic approach with a three-tiered structure comprising physical-to-virtual and virtual-to-virtual communications between layers. This structure forms a comprehensive cyber-twin framework within vehicular ad hoc networks (VANET). Our system is segmented into three layers: data, twinning, and cyber-security, each crucial in ensuring the network’s functionality and resilience, as seen in Fig. 2.
III-A Data Layer
The Data Layer is the foundational segment of our cyber-twin architecture, primarily focusing on data acquisition and transmission. It integrates vehicles and RSUs, collecting critical data for higher-level simulation and security analysis. Its precision in real-time data representation and communication is crucial for maintaining the integrity and effectiveness of the digital twin simulations.
III-B Twinning Layer
This layer is pivotal in creating dynamic digital replicas of VANET’s physical components. It utilizes advanced digital twin modelling and queuing theory to replicate the VANET environment. We have implemented the M/M/m queuing model with a First-In, First-Out strategy for efficient data network simulation. The model evaluates the system’s state by analyzing communication requests against channel availability, providing insights into average waiting times and queue lengths [17].
In this model, channels are likened to servers, and the system’s state is evaluated by comparing the total communication requests against available channels. For the M/M/m queuing model, our key assumptions and parameters are as follows:
  • Vehicles share channels equally within a transmission area.
  • $\lambda_{R}$ as the arrival rate of communication requests.
  • $\mu$, the channel’s service rate per request, time-dependent.
  • $c_{m}$, the total number of channels.
  • $c_{d}$, the total number of vehicles.
These parameters feed into the M/M/m model to determine average wait time and queue length metrics. These parameters are crucial in the M/M/m model for assessing the system’s efficiency and managing channel requests.
We formulate basic equations using a state diagram in Fig. 2. $c$ represent channel states, $m$ represent the number of channels and $d$ denote the number of vehicles with communication requests. When the number of requests $c_{d}$ is less than or equal to the number of available channels $c_{m}$ (i.e., $c_{d}\leq c_{m}$), all requests are accommodated immediately. However, if the requests exceed the available channels (i.e., $c_{d}>c_{m}$), the system processes requests up to the limit of $c_{m}$, and the remaining requests are queued for subsequent processing. There are $c_{m}$ channels so the probability is increasing ($\mu,2\mu,3\mu\;...\;c_{m}\mu$) until reaching $c_{m}{th}$ channel. After the $c_{m}th$ channel, the leaving of the one request stays the same with probability $c_{m}th$.
The utilization factor for this model can be calculated as follows:
$\rho={1-\frac{\lambda_{R}}{c_{m}\mu}}$ (1)
For zero request state:
$P_{0}=\Biggl{[}\sum_{c_{d}=0}^{c_{m}-1}\frac{(c_{m}\rho)^{c_{d}}}{c_{d}\ !}+\frac{(c_{m}\rho)^{c_{m}}}{c_{m}\ !(1-\rho)}\Biggr{]}$ (2)
$P_{c_{d}}$ is the probability of all channels busy in the system:
$P_{c_{d}}=\Biggl{\{}P_{0}\frac{(c_{m}\rho)^{c_{d}}}{c_{d}\ !},\qquad c_{d}\leq c_{m}\\ P_{0}\frac{(c_{m})^{c_{m}}(\rho)^{c_{d}}}{c_{m}\ !},\qquad c_{d}>c_{m}$ (3)
The probability of arriving request has to wait in the queue ( $c_{m}$ request or more in the system):
$P_{Q_{R}}=\sum_{c_{d}=c_{m}}^{\infty}P_{c_{d}}=P_{0}\frac{(c_{m}\rho)^{c_{m}}}{c_{m}\ !(1-\rho)}$ (4)
When the channel is the full state, the average number of requests in the queue:
$N_{Q_{R}}=\frac{P_{Q_{R}}\rho}{(1-\rho)}$ (5)
The amount of time spent in the queue, which is the average waiting time in the queue:
$T_{Q_{R}}=\frac{P_{Q_{R}}\rho}{\lambda_{R}(1-\rho)}$ (6)
The total time in the system, which is the sum of queue time and the service time:
$T=T_{Q_{R}}+\frac{1}{\mu}$ (7)
Lastly, we calculated the average number of channel requests in the system as follows:
$N_{R}={c_{m}}\rho+\frac{P_{Q_{R}}\rho}{1-\rho}$ (8)
Data from the twinning layer is then transferred to the cyber-security layer using a YANG model for robust data transfer.
III-C Cyber-security Layer
The core of our system’s security mechanism is derived from our previous research [18]. It comprises an AutoFS Module and a Multilayer Perceptron, both vital for efficient attack detection.
III-C1 AutoFS Module
This module dynamically selects the most effective Feature Selection (FS) methods under varying network conditions. The system incorporates various techniques such as Recursive Feature Elimination, Backward Feature Elimination, Chi-square, Fisher Score, and ANOVA F-value Selection, adapting to the dynamic nature of network data. Each FS method has a unique approach to feature selection, influencing DDoS detection efficiency. The module identifies the optimal MLP model and FS method based on system performance metrics.
III-C2 Labeling Algorithm
Our approach integrates an ensemble learning strategy, merging the Expectation-Maximization (EM) and K-means algorithms to label data effectively. This combination enhances the precision and efficiency of our data categorization process. This method virtually clusters and labels data, considering the infrequency of the attacks for precise prediction.
III-C3 Multilayer Perceptron (MLP)
Our MLP network is optimized for online learning, adapting to continuously changing network data. Its architectural features are a five-layer architecture with relu and softmax activation functions for effective data classification. It employs online learning, continually updating model weights in response to changing network data, ensuring stable performance in a dynamic network environment.
The proposed system model offers a comprehensive and adaptive approach to vehicular communication, which is crucial for the evolving landscape of smart city infrastructure. Its multi-layered architecture bolsters security and optimizes overall performance, positioning it as a pioneering approach to securing 6G smart cities.
IV Performance Evaluation
To test the performance of our proposed system, we used OMNeT++ version 5.1, INET version 3.6, Veins version 4.7, and SUMO version 0.30.0 as the previous work [6]. Our evaluation focused on two main aspects: the efficiency of the digital twin in RSU resource utilization and the effectiveness of the cyber-twin framework in attack detection.
Firstly, we test the performance of the digital twin to the Road Side Unite (RSU). Therefore, we check the average resource usage of the system with the digital twin and without the digital twin. As seen in Fig. 3, when we employed the cyber-twin in the system, the system utilised the resource more efficiently since we reduced the load of the roadside unit thanks to the digital twin. Integrating the cyber-twin system model resulted in more efficient resource utilization, thereby lightening the RSU’s load.
Figure 3: The average resource usage comparison.
After testing the digital twin effect of the roadside unit resource performance, we experiment with the attack detection performance of the cyber-twin framework. To this end, we used two datasets. The first dataset is the RF Jamming Dataset, which comprises diverse scenarios of RF jamming attacks and interference in VANET [19]. This dataset includes two datasets for two maximum estimated relative speed types. We combined them, adding a new column with the maximum estimated relative speed. The ToN-IoT dataset, specifically designed for assessing the reliability and performance of various AI-driven cybersecurity applications, is utilized in our study [20], since Gad et al. used this dataset for attack detection in VANET. This dataset is particularly tailored for next-generation IoT and industrial IoT scenarios. However, the current version of the ToN-IoT dataset is not VANET-oriented; therefore, we combine the first dataset’s no-attack samples and the ToN-IoT dataset’s attack samples. We selectively used a specified number of samples from these datasets to construct Dataset 2. The details of this selection, including the number of samples from each dataset, are meticulously outlined in Table I.
TABLE I: The Sample Distribution in Dataset 2
Dataset Name (Feature) Number of Samples
RF Jamming Dataset-1
(No Attack Samples)
1000
RF Jamming Dataset-2
(No Attack Samples)
1000
ToN-IoT Network Dataset
(Attack Samples)
400
We compare our solution (PS) with KNN and SVM algorithms since they gave the best result in the following works: [22] and [23], respectively. Fig. 4 and Fig. 5 show the performance comparison of methods for the datasets. As seen in the figures, our solution performs better than the others.
Figure 4: The performance comparison of methods for the dataset 1.
Figure 5: The performance comparison of methods for the dataset 2.
We also investigate the twinning rate performance according to the system’s detection rate and total RAM usage. To this end, we used the dataset 2. We define the twinning rate as follows:
$\gamma=\frac{\tau}{\sigma}*100$ (9)
where $\gamma$ is the twining rate, $\sigma$ is the total number of packages in the roadside units in the data layer, and $\tau$ is the total number of packages taken to the twinning layer from the data layer.
We explored the relationship between the twinning rate, detection rate, and total RAM usage. Fig. 6 depicts the detection rate and total RAM usage comparison according to the twinning rate. According to the results, we defined the optimum twinning rate range between seventy-six and ninety per cent to minimize the total RAM usage while protecting a high attack detection rate. Thus, we established an optimal range that balances resource consumption with high attack detection efficiency by fine-tuning the twinning rate.
Figure 6: The detection rate and total RAM usage comparison according to the twinning rate.
In addition to its robust security and efficiency features, our proposed system significantly contributes to green communication practices in 6G smart city infrastructures. The system’s advanced digital twin technology optimises RSU performance and reduces energy consumption. By efficiently managing data traffic and processing demands, the system minimizes the environmental impact of urban communication networks. This reduction in energy requirements aligns with eco-friendly objectives and diminishes operational costs, making it a sustainable solution for future smart cities. Furthermore, the extended lifespan of hardware components due to reduced strain translates into fewer replacements and maintenance needs, further reinforcing the system’s environmentally conscious design. The cumulative effect of these factors underscores our system’s role in advancing green communication technology, a critical consideration in the development of sustainable urban environments.
V Conclusion
In conclusion, our research has successfully introduced a cyber-twin framework that significantly enhances the security of Road Side Units (RSUs) in Vehicular Ad-hoc NETworks (VANETs). We have developed a robust method for real-time vehicular network monitoring and management by integrating digital twin technology with advanced AI models. This framework adeptly addresses VANETs’ complexities, notably in vehicle-to-infrastructure communications. Our findings demonstrate substantial resource utilization and attack detection improvements, outperforming existing solutions. The cyber-twin framework’s ability to create accurate digital replicas of RSUs has been crucial in effectively detecting and mitigating cyber threats, reinforcing communications’ integrity and reliability. Furthermore, the framework significantly advances green communication practices by enhancing RSU computational efficiency, contributing to energy efficiency and hardware longevity. This aligns with the shift towards sustainable, eco-friendly technologies in intelligent transportation systems and smart cities. Overall, our cyber-twin framework sets a new benchmark in developing secure, sustainable, and intelligent transportation systems, addressing key security challenges while paving the way for a future of safer and greener smart cities.
Acknowledgment
This work is supported by The Scientific and Technological Research Council of Turkey (TUBITAK) 1515 Frontier R&D Laboratories Support Program for BTS Advanced AI Hub: BTS Autonomous Networks and Data Innovation Lab Project 5239903.